OpenStudy (dls):

If the product of three terms in GP is 216 and their sum is 19.Find the GP.

OpenStudy (dls):

\[a*ar*ar ^{2} = 216\] \[a ^{3}r ^{3}\] \[a ^{3}r ^{3}=6^{3}\] ar=6 Am I right till here?

OpenStudy (callisto):

Yes, I think so.

OpenStudy (dls):

so 6+6r+6r^2=19r?

OpenStudy (dls):

idk after this..

OpenStudy (dls):

a+ar+ar^2=19 6/r+6+36/r

OpenStudy (callisto):

Quadratic equation?

OpenStudy (dls):

Is this eq. correct? 42+6r=19r

OpenStudy (dls):

oops its not i guess :/

OpenStudy (dls):

I have this 6(1+r+r^2)=19r

OpenStudy (callisto):

Wow... What's happening :| Just start it again... \[a \times ar \times ar^2 = 216\]\[(ar)^3 = 6^3\]\[ar= 6\]\[r=\frac{6}{a}\] \[a + ar + ar^2= 19\]\[a + 6 + ar^2= 19\]\[a + ar^2= 13\]\[a + a(\frac{6}{a})^2= 13\]Got it so far?

OpenStudy (dls):

Yes I did all this Im stuck here :/ I have this 6(1+r+r^2)=19r

OpenStudy (callisto):

What I did is not the same as what you did. Please check it again. The reason I make r as the subject is because I don't have to change a in the equation.

OpenStudy (dls):

yeah just realized!

OpenStudy (callisto):

Do you understand what I've done then??

OpenStudy (dls):

wont we end up in cubes with this equation? I got this too..

OpenStudy (dls):

oh okay it gets eliminated

OpenStudy (dls):

a^2-13a+36=0

OpenStudy (callisto):

Yup! a quadratic equation!

OpenStudy (dls):

hey i got this before too :o but didn't get factors o.O

OpenStudy (dls):

If I use the formula ill end up in decimal stuff..

OpenStudy (dls):

okay got it thanks !

OpenStudy (callisto):

The answer should be beautiful!